TheSmartOne (thesmartone):

Is my work correct or would I have to do it differently since it's a one sided limit? Find the limit if it exists. If it doesn’t exist, explain why.

TheSmartOne (thesmartone):

\[\lim_{x \rightarrow 3^+}\left(\frac{\sqrt{x}-3}{x-9} \right)\] \[\left(\frac{\sqrt{3}-3}{3-9} \right) = \boxed{\frac{\sqrt{3} - 3}{6}}\]

zepdrix (zepdrix):

Is it maybe a typo? Did they intend for this limit to approach 9 maybe?

zepdrix (zepdrix):

As is, your work looks fine.

TheSmartOne (thesmartone):

Nopee, it's 3 from the right

OpenStudy (zzr0ck3r):

then you did it right

OpenStudy (zzr0ck3r):

Also sure it is not \(\sqrt{x-3}\)?

OpenStudy (zzr0ck3r):

then it would be undefined on the left....

TheSmartOne (thesmartone):

Thanks, next one's coming up. :) The question is http://prntscr.com/cezda4

zepdrix (zepdrix):

Maybe it gets more interesting in b) lol :)

OpenStudy (zzr0ck3r):

weird...sometimes they just want to make sure you can notice the easy ones....I guess.

TheSmartOne (thesmartone):

This question was continuous at 3 so yeah It does get interesting at B, which will come in the next post xD